OpenStudy (anonymous):

solve the system using any algebraic method 4x-y=4 and x+2y=10

OpenStudy (anonymous):

Cross out / cancellation method ( I forgot the actual name of it sorry) cancel out x or y first. So... lets cancel out x , multiply the second equation by the inverse of 4 (negative four) and you get -4x-8y=-40. Set up the first equation above the second 4x-y=4 -4x-8y=-40 the x's cancel out, see? You get -9y=-36 Divide both sides by -9 y = 4 :] Plug four back into one of the *original* equations for y, to solve for x. (I'm new here, hope I explained this well)

OpenStudy (anonymous):

*It's called elimination